# Resource-Efficient Adaptation of Large Language Models for Text Embeddings via Prompt Engineering and Contrastive Fine-tuning

### v1

@@ 4.2 Prompt-Based Shaping of the Embedding Space
Table 5: Zero-shot clustering accuracy across different model sizes, with and without the CCW prompt. The results highlight how model scaling and prompt-based input formatting influence performance, both mean and last-token pooling. The CCW prompt is applied consistently to all prompted variants.
@@
Scaling Llama from 1 billion to 3 billion parameters yields virtually no change in either pooling metric, suggesting that the 1B model already saturates clustering performance, for both zero-shot and with prompting. By contrast, increasing Qwen’s size from 0.6 billion to 4 billion parameters raises last-token accuracy from 30.7 % to 46.3 %, matching Llama’s performance. However, a further increase to 8 billion parameters yields no additional gains in any configuration. Thus, for Qwen, the benefits of scaling plateau before the 8 billion parameter mark, whereas Llama shows minor sensitivity to model size in this task.
4.3 Contrastive Fine-Tuning with Augmented Data
Consistent with prior findings, prompts benefit the last-token variant substantially more than mean pooling. On Llama-3.2-1B, the CCW prompt increases last-token accuracy from 33.2 % to 46.8 %, while mean pooling remains largely unchanged. Qwen3-0.6B follows a similar trend, though with smaller absolute improvements. The EOL prompt even reduces performance for Qwen.
Training LoRA adapters on the human-curated NLI corpus without a prompt improves performance for both models, yet it does not surpasses prompt-based fine-tuning. On Llama, the gap is particularly pronounced — fine-tuning NLI without a prompt achieves 47.8 %, barely outperforming CCW prompting alone at 46.8 %.
The best results are obtained using LLM-generated positive pairs, referred to as “wiki-prompt”. When combined with CCW, last-token accuracy reaches 53.9 % for Llama and 50.9 % for Qwen, representing the overall best scores. Mean pooling also preforms strongly, peaking at 52.7 % for Llama on the same synthetic dataset. For Qwen, mean pooling achieves its best result on NLI data.
Method Mean-pooling
Llama-3.2-1B
Last-token
Llama-3.2-1B
Mean-pooling
Qwen3-0.6B
Last-token
Qwen3-0.6B
Base 45.8 % 33.2 % 39.8 % 28.9 %
EOL 46.1 % 43.2 % 38.3 % 26.9 %
CCW 45.3 % 46.8 % 38.6 % 30.7 %
CCP 43.5 % 32.8 % 38.9 % 29.8 %
Base + nli 49.2 % 47.8 % 46.8 % 47.4 %
CCW + nli 52.5 % 52.2 % 49.2 % 49.6 %
CCP + nli 49.5 % 49.7 % 46.2 % 48.9 %
CCW + wiki-translate 51.2 % 52.2 % 47.1 % 49.4 %
CCP + wiki-translate 47.0 % 45.3 % 44.4 % 43.8 %
CCW + wiki-translate + aug 51.7 % 52.6 % 46.8 % 47.7 %
CCP + wiki-translate + aug 47.2 % 46.5 % 43.4 % 45.6 %
CCW + wiki-prompt 52.7 % 53.9 % 48.8 % 50.9 %
CCP + wiki-prompt 48.4 % 51.8 % 45.3 % 49.0 %
4.3.1 Hyperparameter-Tuning
Table 7 lists the hyperparameters explored during fine-tuning. The final settings used in our experiments are highlighted in bold. We initially trained LoRA adapters on the two attention projections, namely the query and value projections. However, we observed higher scores after extending optimization to include six projections in total. These additional projections also influence the feed-forward pathway and include the output, gate, up, and down projections. This broader parameter coverage is particularly well suited for embedding tasks, as it allows the model to adapt both the attention mechanism and token-mixing components.
Hyperparameter Tested values
Dropout 0.05, 0.1, 0.2, 0.5
Learning rate $1\times 10^{-5}$, $\mathbf{5\times 10^{-5}}$, $5\times 10^{-4}$
Batch size 20, 60, 120, 500
Temperature 1, 0.5, 0.2, 0.1, 0.05
Projection layers 2, 6
All fine-tuning experiments were conducted on a single NVIDIA RTX A5000 GPU with 24 GB of memory. Peak validation performance typically emerged within 0.1 to 0.5 epochs, corresponding to approximately 100 optimization steps or 12,000 examples at the selected batch size of 120. In some cases, performance plateaued later, between 500 and 2,500 steps. To address this, we saved intermediate checkpoints and selected the best-performing model post hoc. In practice, fine-tuning runs completed under 30 minutes, demonstrating the method’s efficiency and suitability for low-resource environments.
4.4 Comparison to Qwen3-Embedding on MTEB Clustering (English)
Qwen3-Embedding-0.6B is explicitly optimized for text embeddings and therefore serves as the most appropriate external baseline for this study. Operating at a comparable scale, it ensures a fair comparison. Our focus lies in efficiently adapting lightweight LLMs, with Qwen3-Embedding providing a state-of-the-art reference within this domain. Although our models are outperformed by the larger Qwen3-Embedding-4B and -8B variants, they achieve competitive results against Qwen3-Embedding-1B. Notably, Qwen3-Embedding-1B surpass nearly all other models on the MTEB, regardless of size, highlighting the strength of this baseline and the competitiveness of our approach. The public MTEB Multilingual v2 leaderboard currently reports Qwen3-Embedding results on only eight English clustering datasets. These are the only tasks for which a direct comparison is possible and are summarized in Table 8.
Dataset/Task Llama_mean Llama_last Qwen3-Emb-0.6B
ArXivHierarchicalClusteringP2P 63.26 % 63.1 % 63.72 %
ArXivHierarchicalClusteringS2S 61.26 % 61.64 % 63.82 %
BigPatentClustering.v2 35.26 % 26.47 % 32.49 %
BiorxivClusteringP2P.v2 42.69 % 42.09 % 47.27 %
MasakhaNEWSClusteringS2S 55.93 % 58.84 % 53.24 %
MedrxivClusteringP2P.v2 37.45 % 38.94 % 42.18 %
SIB200ClusteringS2S 46.41 % 55.52 % 34.13 %
WikiCitiesClustering 85.73 % 80.73 % 86.76 %
AVG 53.5 % 53.42 % 52.95 %
Table 8: Comparison of clustering performance between our best LLaMA models and Qwen3-Embedding-0.6B. Results are reported on the eight English clustering tasks from the MTEB Multilingual v2 leaderboard [24]
Across this subset of tasks, our Llama-3.2-1B checkpoints — fine-tuned with CCW prompt and evaluated using both mean and last-token pooling — achieve average scores of 53.5 % and 53.42 %, respectively, slightly surpassing the 52.95 % reported for Qwen3-Embedding-0.6B. Although Qwen3-Embedding is trained for multi-prompt and multi-task generalization, these results demonstrate that a lightweight, LoRA fine-tuned model can match or exceed its English clustering performance while operating with a fraction of the computational power — less than one hour on a single RTX A5000. Notably, in this restricted setting, the mean pooling outperforms the last-token pooling, reversing the ranking observed on the full English clustering benchmark from MTEB v1.38.30.
4.5 Classification as a Downstream Task
Model Mean-pooling Last-token
Llama-3.2-1B 72.5 % 68.1 %
Llama-3.2-1B + Prompt 73.0 % 70.3 %
Llama-3.2-1B + Prompt + ft 74.2 % 75.4 %
Qwen3-0.6B 70.2 % 63.0 %
Qwen3-0.6B + Prompt 70.3 % 64.1 %
Qwen3-0.6B + Prompt + ft 72.4 % 72.5 %
Table 9: Classification accuracy on the English subset of MTEB using the same configuration as in the clustering experiment. No task-specific prompt tuning or hyperparameter adjustment was applied. A classification-style prompt structurally aligned with the best-performing clustering prompt (CCW) was used.
The observed performance pattern closely mirrors that of the clustering tasks. In the zero-shot setting, mean pooling consistently outperforms last-token pooling, and Llama-3.2-1B maintains a slight advantage over Qwen3-0.6B. Introducing the CCW-style prompt — "This sentence: "[X]" can be classified as:" — yields modest improvements, with again more pronounced for last-token embeddings. Fine-tuning with LoRA using the same hyperparameters as in the clustering experiments increases last-token accuracy to 75.4 % for Llama and 72.5 % for Qwen, establishing this as the best overall configuration. Notably, these improvements are obtained without classification-specific prompt tuning or hyperparameter optimization, highlighting the transferability of the pipeline across task types.
We further evaluate a key advantage of our method — its ability to incorporate domain-specific knowledge through targeted fine-tuning. To this end, we fine-tune Llama-3.2-1B on two MTEB classification datasets with labeled training splits: EmotionClassification and ToxicConversationsClassification[10]. Based on prior results showing its superiority in fine-tuned settings, we use last-token pooling exclusively. For each task, we construct a domain-specific contrastive training set by prompting an LLM to generate paraphrases for each sentence in the task’s training set, thereby creating positive pairs. This mirrors the "wiki-prompt" approach, but substitutes Wikipedia with in-domain text. Fine-tuning on these smaller classification datasets is even faster than in the clustering setting. As before, less than a full epoch is required, with performance typically peaking after approximately 100 optimization steps.
Method Emotion Toxic
Last 29.26 % 63.53 %
Last + prompt 42.81 % 63.19 %
Fine-tuned (general) 45.33 % 67.38 %
Fine-tuned (task) 49.78 % 68.93 %
Table 10: Classification accuracy on EmotionClassification and ToxicConversationsClassification. “Last” denotes the last-token embedding without a prompt; “+ prompt” adds a classification-style prompt. “Fine-tuned (general)” refers to LoRA training on the wiki-prompt corpus, while “Fine-tuned (task)” uses LLM-generated positive pairs derived from each task’s labeled training set.
Table 10 illustrates a consistent pattern. The classification prompt leads to a substantial improvement on the Emotion task but yields minimal gains for Toxicity. Fine-tuning on the general wiki-prompt corpus enhances performance across both tasks. However, fine-tuning on the task-specific data achieves the highest scores, improving Emotion by 4.5 % and Toxicity by 1.6  % over the general model. While the extent of improvement varies depending on the domain, the addition of task-relevant supervision consistently results in further gains.
4.6 Analysis of Attention
To investigate how LoRA fine-tuning alters the model’s internal focus, we follow a method similar to that of Zhang et al. [1]. Attention maps were generated for the probe sentence "a man is driving a car" and validated using additional sentences sampled from MTEB. The analysis examines the outgoing attention distribution of the final token in Llama-3.2-1B with the CCW prompt, comparing three settings illustrated in Figure 1: panel (a) shows the zero-shot model, panel (b) presents the best checkpoint after approximately 100 optimization steps, and panel (c) displays an over-trained checkpoint after 2500 steps.
In all cases, the final token directs most of its attention to itself, a well-known behavior in causal transformers that preserves the accumulated context through residual connections. The key difference lies in the distribution of the remaining attention. After moderate fine-tuning, attention shift away from the prompt tokens toward the content words man, driving, and car, emphasizing the sentence’s subject, predicate, and object — the elements most informative for downstream semantic tasks. With extended training, the focused attention dissipates; attention becomes more uniformly distributed, and model performance declines accordingly. These observations suggest that an appropriately timed fine-tuning phase encourages the model to encode task-relevant semantics into the final hidden state, while over-training diffuses this information and diminishes embedding quality.
Figure 1: Average attention received by each input token from the final token in the last transformer layer of LLaMA-3.2-1B, prompted with "This sentence: ‘a man is driving a car’ belongs to the following cluster:". Bars indicate values for at three stages of training: prompt-only, after 100 fine-tuning steps, and after 20,000 steps. Fine-tuning initially shifts attention to semantically meaningful tokens (“man”, “driving”, “car”), while prolonged training leads to a more uniform distribution across the sequence.
5 Conclusion
This work focuses on turning pretrained LLMs into strong sentence-embedding models. We show that parameter-efficient fine-tuning, such as LoRA steered by a well-chosen prompt, can achieve high-quality embeddings. Additionally, sentence-level data augmentation like synthetic positive pairs enables promising contrastive self-supervised fine-tuning without manually labeled data. We raised Llama-3.2-1B and Qwen3-0.6B to state-of-the-art performance on the English clustering subset of MTEB, surpassing a dedicated embedding baseline in under an hour on a single RTX A5000. Attention maps reveal that our fine-tuning paradigm leads to a more interpretable adaptation to downstream tasks.
Future work will explore task-agnostic prompt formats, extend the framework to multilingual and cross-modal applications, and investigate quantization techniques to further reduce resource requirements.
Acknowledgements

### v2

@@ B.1 Zero-Shot Comparison by Model Size
Table 10: Zero-shot clustering score across different model sizes, with and without the CCW prompt. The results highlight how model scaling and prompt-based input formatting influence performance, both mean and last-token pooling. The CCW prompt is consistently applied.
@@
Scaling Llama from 1 billion to 3 billion parameters yields virtually no change in either pooling metric, suggesting that the Llama-3.2-1B model already saturates clustering performance, for both zero-shot and with prompting. By contrast, increasing Qwen’s size from 0.6 billion to 4 billion parameters raises last-token accuracy from 30.7 % to 46.3 %, matching Llama’s performance. However, a further increase to 8 billion parameters yields no additional gains in any configuration. Thus, for Qwen, the benefits of scaling plateau before the 8 billion parameter mark, whereas Llama shows minor sensitivity to model size in this task.
B.2 Detailed Performance and Standard Deviation of the Best Model
Dataset/Task Llama_last
ArXivHierarchicalClusteringP2P 62.71 $\pm$ 02.34 %
ArXivHierarchicalClusteringS2S 63.00 $\pm$ 04.13 %
BigPatentClustering.v2 32.11 $\pm$ 03.44 %
BiorxivClusteringP2P.v2 42.35 $\pm$ 00.57 %
BiorxivClusteringS2S.v2 38.80 $\pm$ 00.56 %
BuiltBenchClusteringP2P 58.22 $\pm$ 12.50 %
BuiltBenchClusteringS2S 50.69 $\pm$ 10.74 %
ClusTREC-Covid 84.89 $\pm$ 00.45 %
MasakhaNEWSClusteringP2P 64.15 $\pm$ 23.64 %
MasakhaNEWSClusteringS2S 59.09 $\pm$ 23.70 %
MedrxivClusteringP2P.v2 39.14 $\pm$ 00.78 %
MedrxivClusteringS2S.v2 37.21 $\pm$ 00.68 %
RedditClustering.v2 58.91 $\pm$ 00.85 %
RedditClusteringP2P.v2 59.72 $\pm$ 00.37 %
SIB200ClusteringS2S 54.67 $\pm$ 05.04 %
StackExchangeClustering.v2 65.67 $\pm$ 00.75 %
StackExchangeClusteringP2P.v2 43.73 $\pm$ 00.43 %
TwentyNewsgroupsClustering.v2 53.03 $\pm$ 01.50 %
WikiCitiesClustering 80.68 $\pm$ 00.00 %
WikipediaChemistryTopicsClustering 67.47 $\pm$ 00.00 %
WikipediaSpecialtiesInChemistryCl. 28.08 $\pm$ 00.00 %
AVG 54.49 %
Table 11: Clustering accuracy of Llama_last across multiple MTEB tasks. Results show the v_measure and v_measure_std reported by running MTEB using mteb.get_tasks(task_types="Clustering", languages=["eng"]).
In Table 11 we report per-dataset v-measure and standard deviation for all 21 English clustering tasks in MTEB v1.38.30. The ± values are not derived from repeated experimental runs, but from the variability induced by MTEB’s own evaluation procedure. Specifically, for clustering tasks MTEB fixes the dataset splits and embeddings, but the downstream clustering evaluation involves multiple runs (e.g., different initializations or folds), and the reported standard deviation quantifies this within-task variation. It should therefore be interpreted as a measure of how sensitive the evaluation metric is to these internal clustering runs, not as a statistical confidence interval. The benchmark suite spans diverse domains, including scientific abstracts (ArXiv, BioRxiv, MedRxiv), news (MasakhaNEWS), online forums (Reddit, StackExchange), and encyclopedic text (Wikipedia, WikiCities). It further distinguishes between task granularities, such as Sentence-to-Sentence (S2S) tasks, which use short text segments like titles or individual sentences, and Passage-to-Passage (P2P) tasks, which rely on longer text spans such as abstracts or full passages [10]. This coverage ensures that the results reflect robustness across heterogeneous text types. The macro-average of 54.49 % indicates stable overall performance, with the strongest results on well-structured corpora such as ClusTREC-Covid and WikiCities. When comparing tasks available in both configurations, P2P generally outperforms S2S, as observed for BioRxiv, BuiltBench, MasakhaNEWS, MedRxiv, and Reddit. However, this trend is not universal: ArXiv shows nearly identical performance for both settings, while StackExchange exhibits a strong advantage for S2S over P2P. These results indicate that the optimal task granularity is dataset-dependent, with P2P often providing richer contextual signals but S2S remaining competitive or superior when concise text segments are more informative. Different tasks reach their best performance after varying numbers of training iterations. Because evaluation was conducted jointly across all 21 tasks, checkpoints were selected based on overall performance rather than task-specific peaks. As a result, the chosen checkpoints may not be optimal for every individual dataset. All results were obtained with the official MTEB evaluation pipeline [10], ensuring direct comparability with prior work.
B.3 Analysis of Attention
In this section, we have chosen randomly several additional example sentences for a visual attention analysis. Each plot shows the average attention received by every input token from the final token in the last transformer layer of Llama-3.2-1B. For the fine-tuned condition, we use the best-performing model. Bars indicate attention at three stages of training: prompt-only, after 100 fine-tuning steps, and after 20,000 steps. Across all examples, the pattern is consistent: short fine-tuning shifts attention from prompt tokens toward the actual sentence being encoded, improving over the prompt-only baseline, whereas prolonged fine-tuning results in a more uniform distribution. The sentences were chosen to be short for readability. In addition to examples aligned with [1], we include samples from the EmotionClassification dataset [10] as well as other simple sentences.
Figure 3: Average attention received by each input token from the final token in the last layer of Llama-3.2-1B with the CCW prompt. Shown for two EmotionClassification sentences [10]. Each plot reports three stages: prompt-only, after 100 fine-tuning steps, and after 20,000 steps.
Figure 4: Average attention received by each input token from the final token in the last layer of Llama-3.2-1B with the CCW prompt. Shown for two random short sentences: “I feel tired after running” and “She read a short story.” Each plot reports three stages: prompt-only, after 100 fine-tuning steps, and after 20,000 steps.
Figures 3 and 4 show how fine-tuning changes attention patterns across four sentences. For the EmotionClassification examples, attention shifts from prompt tokens toward emotionally salient words such as embarrassed, worthless, and struggling. In the shorter generic sentences, the model similarly learns to emphasize core content words like tired, running, read, and story. Extended training eventually redistributes attention more evenly while still retaining stronger focus on semantically central tokens.
Overall, the zero-shot setting already reveals that semantically relevant tokens tend to attract higher attention, indicating that the model encodes some notion of relevance without task-specific adaptation. Fine-tuning maintains this advantage while shifting focus away from prompt tokens toward the sentence to be embedded. The combination of prompt usage, which guides the model to store information in the final token, and fine-tuning, which encourages greater focus on semantic content, indicates that the model learns to produce more discriminative embeddings.